Derive an expression for the electric field at a point :
(i) along the axis of a dipole, and
(ii) on the perpendicular bisector of the
dipole axis
1
Expert's answer
2019-05-24T09:55:53-0400

(i)





"E=E_1+E_2 (1)"

where


"E_1=\frac { 1}{4\pi ɛ_0}\frac { q}{(r-a) ^2}"


"E_2=\frac { 1}{4\pi ɛ_0}\frac { -q}{(r+a) ^2}"

Using (1) we got:


"E=\frac { 1}{4\pi ɛ_0} 4aq \frac { r}{(r^2-a^2) ^2} (2)"

The dipole moment (p) is given by formula


"p=2qa (3)"

We put (3) in (2):


"E=\frac { 1}{2\pi ɛ_0} \frac { pr}{(r^2-a^2) ^2} (4)"

Consider the far field case (i.e., the case where r >> a)

Then one can see that r2 >> a2, which allows us to simplify the equation in the far field.


"E=\frac { 1}{2\pi ɛ_0} \frac { p}{r^3}"

(ii)






"E=2E_1 \ cos θ (1)"

where


"E_1=\frac { 1}{4\pi ɛ_0} \frac { q}{a ^2+x^2}"

"\ cos θ=\frac { a}{\sqrt { a ^2+x^2}}"

Using (1) we got:


"E=\frac { 1}{4\pi ɛ_0} \frac { 2qa}{ \sqrt[3]{ a ^2+x^2}}"

By inspection we can conclude that when x>>a then this approaches


"E=\frac { 1}{4\pi ɛ_0} \frac { 2qa}{ x^3}"